A pizza parlor has a choice of 9 toppings for its pizzas. From these 9 toppings, how many different 2 topping pizzas can be ordered? Assume that the order in which the toppings are listed does not matter.
Answer by stanbon(75887)

A pizza parlor has a choice of 9 toppings for its pizzas. From these 9 toppings, how many different 2 topping pizzas can be ordered? Assume that the order in which the toppings are listed does not matter.
-----
Ans: 9C2 = (9*8)/(1*2) = 9*4 = 36
---------------
Cheers,
Stan H.
-----------